Color code: (critical, news during the meeting: green, news from this week: blue, news from last week: purple, no news: black)

Problem with EoS  and START/STOP/START:

Full system test:

Global calibration status:

Failures reading CCDB objects at high rate.

Issues currently lacking manpower, waiting for a volunteer:

Speeding up startup of workflow on the EPN:

EPN Scheduler

Important framework features to follow up:

Minor open points:

Non-critical QC tasks:

Reconstruction performance:

Topology generation:

TPC DLBZS / SYNTHETIC runs

O2 versions at P2:

Other SRC items:

 

QC / Monitoring / InfoLogger updates:

Testing Pb-Pb workflow at P2: